# Microscopic model for aging of biocondensates

### v2

@@ Results
This non-linear differential equation cannot be solved analytically; however, we show in the Supplementary Material that for a large systems ($N$ and $\bar{n}$ large) $\bar{n}\propto\log(t)$. This logarithmic scaling can be used together with the dominant factor of Eq. (8): $\tau(\bar{n})\propto e^{\bar{n}}$. Which gives $\tau(t)\propto t^{\gamma}$, where $\gamma$ is a constant that depends on $\tau_{\text{exch}}$. This power-law qualitatively agrees with the inset of Fig. 4 (b), which confirms the relation between the relaxation time of the ISF, and the dissolution time of growing clusters. Here, we focused on the dominant mathematical factors, and neglected many processes that could provide a more quantitative agreement with the simulation. This choice is motivated by the initial strong simplification of our sticker and spacer model. And we believe that only qualitative results will be robust to the addition of real microscopic details.
Figure 4: (a) average cluster size evolution, clusters are built using a hierarchical algorithm that use a minimum distance linkage criteria. The distance used is the average sticker distance computed using the pair correlation function of Eq. (5), in our case $\bar{d}=3.6$. (b) Modulus, of the intermediate scattering function for $|\bm{k}|=1/\bar{r}=0.37$, where $\bar{r}$ is the average distance between stickers. The stretched exponential fits are in good agreement, and the corresponding $\tau$ values obtained are plotted in the inset. The value of $\alpha$ does not evolve significantly over time, and remains $\approx 0.7$.
@@ Discussion
We have introduced a minimal model to describe the dynamics of protein or RNA/DNA-protein condensates. By utilizing the general sticker and spacer framework, we examined the long-term structural evolution of these systems. Our investigation primarily focused on two general effects: the entropic cost associated with binding and the increasing difficulty of the system to rearrange upon equilibration. Our model reveals two distinct mechanisms: the emergence of a Casimir attractive force between stickers mediated by spacers and the clustering of stickers leading to glassy relaxation. The first mechanism is reminiscent of bridging-induced phase separation, where RNA or DNA-protein binding is driven by a similar entropic force, triggering condensation [3, 9]. In our framework, this entropic attractive force appears for high binding energy, as the simultaneous unbinding of multiple stickers would disrupt the attractive interaction This approximation might not be universally valid for all biological systems. in fact, we anticipate that in many cases, the Casimir force will not be the primary driver of structural reorganization. Nonetheless, in our model, this attractive force is only useful to drive sticker clustering, which is ultimately responsible for glassy relaxation. This clustering can also be driven by system-specific mechanisms, such as three-body stabilizing interactions or multiple $\beta$-sheet stacking, and has been observed in experimental systems [25, 15, 32]. In fact, such clustering is also a general behavior of random walkers in disordered media, commonly known as localization in this context [33]. For these reasons, we are confident that the clustering of stickers described here is a fairly universal feature, from which we have only highlighted the entropic minimization effect.
@@
Our model shares similarities with previously studied models of random walker with trap [34, 35], which reproduce key features, such as glass transition, subdiffusion, or localization, reminiscent of the phenomenology observed here. However, these models use a prior distribution of binding energies (e.g., exponential), aging occurs through the exploration of deeper and deeper traps by the random walker. On the one hand, real proteins exhibit a distribution of interaction energy, which is sufficient to induce complex relaxation of a single protein [36]. On the other hand, proteins’ interactions –electrostatic, hydrophobic, or hydrogen bonding– are all of the same order of magnitude ($\approx 10k_{B}T$), which should lead to a fairly narrow distribution of interaction energy. For this reason, intrinsic protein complexity does not justify a priori the existence of ever-deepening traps needed for long-time aging. Our model offers a rationale for considering a broad distribution of energetic interactions, as proposed in previous studies [19, 37]. By focusing on the exponential part of Eq. (8), we find that the dissolution time of clusters is similar to a Kramer escape time for a trap of depth $\bar{n}$ time larger than an isolated sticker. Such mechanism is known for enhancing small heterogeneities of stickers, leading to complex viscoelastic relaxation [38].
@@ Acknowledgements.
We thank Lara Koehler for her critical reading of the manuscript. This work was supported by the Swiss National Science Foundation under grant CRSII5 193740.

### v3

@@ Results
This non-linear differential equation cannot be solved analytically; however, we show in the Supplementary Material that for a large systems ($1\ll\bar{n}\ll N$ ) : $\bar{n}\propto\log(t)$. This logarithmic scaling can be used together with the dominant factor of Eq. (8): $\tau(\bar{n})\propto e^{\bar{n}}$. Which gives $\tau(t)\propto t^{\gamma}$, where $\gamma$ is a constant that depends on $\tau_{\text{exch}}$. This power-law qualitatively agrees with the inset of Fig. 4 (b), which confirms the relation between the relaxation time of the ISF, and the dissolution time of growing clusters. Here, we focused on the dominant mathematical factors, and neglected many processes that could provide a more quantitative agreement with the simulation. This choice is motivated by the initial strong simplification of our sticker and spacer model. And we believe that only such qualitative results will be robust to the addition of real microscopic details.
Figure 4: (a) Average cluster size evolution, clusters are built using a hierarchical algorithm that use a minimum distance linkage criteria. The distance used is the average sticker distance computed using the pair correlation function of Eq. (5), in our case $\bar{d}=3.6$. (b) Modulus, of the intermediate scattering function for $|\bm{k}|=1/\bar{r}=0.37$, where $\bar{r}$ is the average distance between stickers. The stretched exponential fits are in good agreement, and the corresponding $\tau$ values obtained are plotted in the inset. The value of $\alpha$ does not evolve significantly over time, and remains $\approx 0.7$. Notice the abrut decay for small $t_{\text{lag}}$ that corresponds to the initial binding regime.
Viscoelastic response
We now use our measurements of the intermediate scattering function (ISF) to estimate the long-time evolution of the dynamic relaxation modulus of our system. The mechanical relaxation of a chemically crosslinked polymer gel occurs through the unbinding of stickers. When a loaded bond breaks, the stress stored in that cross-link is relieved. If the unbound sticker has sufficient mobility, future rebinding occurs without reintroducing stress into the network. In our model, bound stickers remain fixed, while unbound stickers diffuse. Consistent with our previous assumptions, we assume that unbound stickers diffuse rapidly compared to the relaxation dynamics of the polymer network.
Based on these principles, we derive in the Supplementary Material a relation connecting the ISF to the dynamic modulus for wave vector $k=1/\langle r\rangle$:
$G(\omega)=\frac{1}{\omega j}\int_{0}^{+\infty}e^{-j\omega t}I(t,t_{\text{lag}}),\text{d}t,$ (10)
where $\Re[G(\omega)]=G^{\prime}(\omega)$ is the storage modulus, and $\Im[G(\omega)]=G^{\prime\prime}(\omega)$ is the loss modulus at angular frequency $\omega$.
The results are shown in Fig. 5 and are consistent with previous experimental [13] and simulation results [34]. As expected from the growing relaxation timescales of the ISF, the crossover frequency—where the storage modulus exceeds the loss modulus, marking the transition from liquid-like to solid-like behavior-shifts to lower values as the lag time increases. This confirms that the material becomes progressively more solid-like at long times.
To aid interpretation, we also provide reference scaling laws characteristic of Maxwell viscoelastic behavior. Although the measured curves resemble those of a Maxwell material, the scalings differ slightly. This deviation arises because the ISF follows a stretched exponential form rather than a single exponential. However, the stretching exponent $\alpha$ remains close to 1, leading to a viscoelastic response that closely approximates Maxwellian behavior.
Figure 5: Estimated storage ($G^{\prime}$) and loss ($G^{\prime\prime}$) modulus, obtained from Eq. 10, using ISF measured after different lag-time. Slopes characteristics of Maxwell viscoelastic gel are displayed as reference.
@@ Discussion
We have introduced a minimal model to describe the dynamics of protein or RNA/DNA-protein condensates. By utilizing the general sticker and spacer framework, we examined the long-time structural evolution of these systems. Our investigation primarily focused on two general effects: the entropic cost associated with binding and the increasing difficulty of the system to rearrange upon equilibration. Our model reveals two distinct mechanisms: the emergence of a Casimir attractive force between stickers mediated by spacers and the clustering of stickers leading to glassy relaxation. The first mechanism is reminiscent of bridging-induced phase separation, where RNA or DNA-protein binding is driven by a similar entropic force, triggering condensation [3, 9]. In our framework, this entropic attractive force appears for high binding energy, as the simultaneous unbinding of multiple stickers would disrupt the attractive interaction. This approximation might not be universally valid for all biological systems. in fact, we anticipate that in many cases, the Casimir force will not be the primary driver of structural reorganization. Nonetheless, in our model, this attractive force is only useful to drive sticker clustering, which is ultimately responsible for glassy relaxation. This clustering can also be driven by system-specific mechanisms, such as three-body stabilizing interactions or multiple $\beta$-sheet stacking, and has been observed in experimental systems [27, 15, 22]. In fact, such clustering is also a general behavior of random walkers in disordered media, commonly known as localization in this context [35]. For these reasons, we are confident that the clustering of stickers described here is a fairly universal feature, from which we have only highlighted the entropic minimization effect.
@@
Our model shares similarities with previously studied models of random walker with trap [36, 37], which reproduce key features, such as glass transition, subdiffusion, or localization, reminiscent of the phenomenology observed here. However, these models use a prior distribution of binding energies (e.g., exponential), aging occurs through the exploration of deeper and deeper traps by the random walker. On the one hand, real proteins exhibit a distribution of interaction energy, which is sufficient to induce complex relaxation of a single protein [38]. On the other hand, proteins’ interactions, electrostatic, hydrophobic, or hydrogen bonding, are all of the same order of magnitude ($\approx 10k_{B}T$), which should lead to a fairly narrow distribution of interaction energy. For this reason, intrinsic protein complexity does not justify a priori the existence of ever-deepening traps needed for long-time aging. Our model offers a rationale for considering a broad distribution of energetic interactions, as proposed in previous studies [19, 39]. By focusing on the exponential part of Eq. (8), we find that the dissolution time of clusters is similar to a Kramer escape time for a trap of depth $\bar{n}$ time larger than an isolated sticker. Such mechanism is known for enhancing small heterogeneities of stickers, leading to complex viscoelastic relaxation [40].
@@ Acknowledgements.
We thank Ned Wingreen for insightful discussions that contributed to improve this work and Lara Koehler for her thorough and critical reading of the manuscript. This research was supported by the Swiss National Science Foundation (grant CRSII5 193740).
Data Availability
All the data presented in this article come from c++ and python simulation. The source codes are available at : https://github.com/HugoLeRoy94/cpp_file_aging_condensate.git and https://github.com/HugoLeRoy94/Parallel_gillespie.git